# What’s Mine becomes Yours: Defining, Annotating and Detecting Context-Dependent Paraphrases in News Interview Dialogs

### v1

@@ Abstract
Best practices for high conflict conversations like counseling or customer support almost always include recommendations to paraphrase the previous speaker. Although paraphrase classification has received widespread attention in NLP, paraphrases are usually considered independent from context, and common models and datasets are not applicable to dialog settings. In this work, we investigate paraphrases in dialog (e.g., Speaker 1: “That book is mine.” becomes Speaker 2: “That book is yours.”). We provide an operationalization of context-dependent paraphrases, and develop a training for crowd-workers to classify paraphrases in dialog. We introduce a dataset with utterance pairs from NPR and CNN news interviews annotated for context-dependent paraphrases. To enable analyses on label variation, the dataset contains 5,581 annotations on 600 utterance pairs. We present promising results with in-context learning and with token classification models for automatic paraphrase detection in dialog.
@@ 1 Introduction
Dataset Acc. MV K’s $\alpha$ Shortened Example Vote
BALANCED
0.71
0.32
@@
G: […] Maybe the money will help. H: It can’t hurt, let’s put it that way.
9/20
RANDOM
0.72
0.23
@@
G: So both parties agree that we need to stop horrific acts of violence against animals. But everyone is standing behind this. It is time to stop horrific acts of brutality on animals. Host: Britain’s Queen Elizabeth’s senior dresser writes "If her majesty is due to attend an engagement in particularly cold weather from 2019 onwards fake fur will be used to make sure she stays warm." it’s a very stark example of a monarch following public opinion in the U.K. which is moving away from fur and it very much embraces prevention of cruelty to the animals.
7/15
PARA
0.65
0.19
@@
G: […] it could be programmed in. But again, you’d have to set that up as part of your flight plan. H: So you’d have to say I’m going to drop to 5,000 feet, then go back up to 35,000 feet, and you would have had to have done that at the beginning.
8/15
@@
Table 1: Agreement Scores as an Indicator of Plausible Variation. For each dataset, we display (Acc. MV) the “accuracy” with the majority vote which is the mean overlap of a rater’s classification with the majority vote classification excluding the current rater and (K’s $\alpha$) Krippendorff (1980)’s alpha for the binary classifications by all raters over all pairs. The relatively low K’s $\alpha$ scores can be explained by pairs where either label is plausible. We display such an example for each dataset with the share of annotators classifying it it as a paraphrase (Vote).
@@
Fortunately, in NLP, paraphrases have received wide-spread attention: Researchers have created numerous paraphrase datasets (Dolan and Brockett, 2005; Zhang et al., 2019; Dong et al., 2021; Kanerva et al., 2023), developed methods to automatically identify paraphrases (Zhang et al., 2019; Wei et al., 2022a; Zhou et al., 2022), and used paraphrase datasets to train semantic sentence representations (Reimers and Gurevych, 2019; Gao et al., 2021) and to benchmark LLMs (Wang et al., 2018; bench authors, 2023). However, most previous work (1) has focused on context-independent paraphrases, i.e., texts that are semantically equivalent independent from the given context, and has not investigated the automatic detection of paraphrases in dialog specifically, (2) has classified paraphrases at the level of full texts even though paraphrases often only occur in portions of larger texts (see also Figure 1), (3) uses a small number of 1–3 annotations per paraphrase pair (Dolan and Brockett, 2005; Kanerva et al., 2023) and (4) only annotate text pairs that are “likely” to include paraphrases using heuristics such as lexical similarity (Dolan and Brockett, 2005), although, especially for the dialog setting, we can not expect lexical similarity to be high for all or even most paraphrase pairs (e.g., the pair in Figure 1 only overlaps in two words).
@@
We address all four limitations with this work. First, we are, to the best of our knowledge, the first to focus on operationalizing, annotating and automatically detecting context-dependent paraphrases in dialog. Dialog is a setting that is uniquely sensitive to context (Grice, 1957, 1975; Davis, 2002), e.g., “doing this for a while now” and “20th time […] as a television commentator” in Figure 1 are not generally semantically equivalent. Second, instead of classifying whether two texts are paraphrases of each other, we focus on classifying whether a text paraphrases (part) of a previous text, and identifying the text spans that constitute the paraphrase pair (e.g., Figure 1). Third, we collect a larger number of annotations of up to 21 per item in line with typical efforts to address plausible human label variation (Nie et al., 2020; Sap et al., 2022). Even though context-dependent paraphrase identification in dialog might at first seem straight forward with a clear ground truth, similar to other “objective” tasks in NLP (Uma et al., 2021), human annotators (plausibly) disagree on labels (Dolan and Brockett, 2005; Kanerva et al., 2023). For example, consider the first text pair in Table 1. “[The money] can’t hurt” can be interpreted in at least two different ways: as a statement with approximately the same meaning as “the money will help” or as an opposing statement meaning the money actually won’t help but at least “It can’t hurt” either. Fourth, instead of using heuristics to select text pairs for annotations, we choose a setting where paraphrases are relatively likely to occur: transcripts of NPR and CNN news interviews (Zhu et al., 2021) since in (news) interviews paraphrasing or more generally active listening is encouraged (Clayman and Heritage, 2002; Hight and Smyth, 2002; Sedorkin et al., 2023).
@@
In short, we operationalize context-dependent paraphrases in dialog with a definition and an iteratively developed hands-on training for annotators. Then, annotators classify paraphrases and identify the spans of text that constitute the paraphrase. We release a dataset with 5,581 annotations on 600 utterance pairs from NPR and CNN news interviews. We use in-context learning (ICL) with popular generative models like Llama 2 or GPT-4 and fine-tune a DeBERTa token classifier to detect paraphrases in dialog. We reach promising results of F1 scores from 0.66 to 0.81. Generative models perform better at classification, while the token classifier provides better text spans. We hope to ultimately advance the reliable detection of paraphrases in dialog. Code and data will be made publicly available [^1] GITHUB LINK WILL BE ADDED AFTER BLIND REVIEW for research purposes in line with Zhu et al. (2021).
@@
Approximate Contextual Equivalence $\subseteq$ CP
G: I’m like, "Fortnite", what is that? I don’t even know what it is – H: So, you weren’t even familiar?
@@
G: There are militant groups out there firing against the military. H: Why did the army decide today to move in and clear out the camp?
@@ 2 Related Work
Paraphrase taxonomies commonly go beyond binary paraphrase classifications and make more fine-grained distinctions between paraphrase types (Bhagat and Hovy, 2013; Vila et al., 2014; Kovatchev et al., 2018; Kanerva et al., 2023). However, context has been sparsely discussed in paraphrase taxonomies, even though considering the context can reverse semantic relations (Shwartz and Dagan, 2016). Even when context is considered, it is not the focus (Bhagat and Hovy, 2013), annotations are not provided (Vila et al., 2014), it is not investigated for dialogs (Kovatchev et al., 2018; Kanerva et al., 2023), or definitions include pairs that fall outside our understanding of a context-dependent paraphrase (Vila et al., 2014). [^2] For example, Vila et al. (2014) encompass text pairs with the same communicative intention like “Close the door please” and “There is air flow”.
@@ 3 Context-Dependent Paraphrases in Dialog
We define a context-dependent paraphrase as two text excerpts that are approximately equivalent in meaning in a given situation but not necessarily in all non-absurd situations.333definition combines elements from Kanerva et al. (2021) and Bhagat and Hovy (2013) For example, consider the first exchange in Table 2. In this situation, “I” uttered by the first speaker and “You” uttered by the second speaker are clearly signifying the same person. However, if uttered by the same speaker “I” and “you” probably do not signify the same person. The text pair in Table 2 is thus equivalent in at least one but not in all non-absurd situations. The text excerpts forming context-dependent paraphrases do not have to be complete utterances. In many cases they are portions of utterances, see highlights in Figure 1. Note that in dialog, the second speaker should rephrase part of the first speaker’s point in the given situation (context condition) and not just talk about something semantically related (equivalence condition).
@@
We specifically exclude common cases of disagreements between annotators444 derived from pilot studies, see also App. C.1 and specifically App. Table 14 that we consider not to be context-dependent paraphrases in dialog, see Table 3. First, we exclude text spans that might be considered approximately equivalent when they are looked at in isolation but do not represent a paraphrase of the guest’s point in the given situation (e.g., “the military” and “the army” in Table 3). Second, we exclude text pairs that diverge too much from the original meaning when the second speaker adds conclusions, inferences or new facts. In an interview setting, journalists make use of different question types and communication strategies relating to their agenda (Clayman and Heritage, 2002) that can sometimes seem like paraphrases. For example in Table 3, the host’s question “So, you …?” could be read as a paraphrase with the goal of checking understanding with the guest. However, it is more likely to be a declarative conclusion that goes beyond what the guest said.
@@ 4 Dataset
Generally, people do not paraphrase each other in every conversation. Therefore, we focus on the news interview setting, because paraphrasing, or more generally active listening, is a common practice for journalists (Clayman and Heritage, 2002; Hight and Smyth, 2002; Sedorkin et al., 2023). We therefore also only consider whether the journalist (the interview host) paraphrases the interview guest and not the other way around. We use Zhu et al. (2021)’s MediaSum corpus which consists of over 450K news interview transcripts and their summaries from 1999–2019 NPR and 2000–2020 CNN interviews. [^5] Released for research purpose, see https://github.com/zcgzcgzcg1/MediaSum?tab=readme-ov-file.
@@ 4.2 Data Samples for Annotation
Even though paraphrases are relatively likely in the news interview setting, most randomly sampled text pairs still do not include paraphrases. To distribute annotation resources to text pairs that are likely to be paraphrase, previous work usually selects pairs based on heuristics like textual similarity features, e.g., word overlap, edit distance, or semantic similarity (Dolan and Brockett, 2005; Su and Yan, 2017; Dong et al., 2021). However, these approaches are systematically biased towards selecting more obvious, often lexically similar text pairs, possibly excluding many context-dependent paraphrases. For example, the guest and host utterance in Figure 1 have varying lengths, only overlap in three words and have a semantic similarity score of only 0.13666 using cosine-similarity and encodings from https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Similar to Kanerva et al. (2023), we instead use a manual selection of promising text pairs for annotation: We (1) randomly sample a set of text pairs and (2) manually classify at each of them to (3) select three sets of text pairs that vary in their paraphrase distribution for the more resource-intensive crowd-sourced annotations: the RANDOM, BALANCED and PARA set.
@@
Lead Author Annotation. We shuffle and uniformly sample 1,304 interviews. For each interview, we sample a maximum of 5 consecutive (guest, host)-pairs. To select promising paraphrase candidates, the lead author then manually classifies all 4,450 text pairs as paraphrases vs. non-paraphrases (see App. B.2 for details).777 After experimenting with crowd-workers, having a first pass for selection done by one of our team seemed the best considering cost-performance trade-offs. In total, about 14.9% of the sampled text pairs are classified as paraphrases by the lead author. On a random set of 100 (guest, host)-pairs (RANDOM), we later compare the lead author’s classifications with the crowd-sourced paraphrase classifications (see App. 12). 89% of the lead author’s classifications are the same as the crowd majority. Note that the lead author’s classifications do not affect the quality of the annotations released with the dataset but only the text pairs that are selected for annotation. However, using lead author annotations instead of lexical level heuristics should increase paraphrase diversity in the released dataset beyond high lexical similarity pairs.
@@ 5.1 Annotation Task
Given a (guest, host) utterance pair, the task annotators perform is (1) classifying whether the host is paraphrasing any part of the guest’s utterance and, if so (2) highlighting the paraphrase in the guest and host utterance. This results in data points like the one in Figure 1. Note that in contrast to previous work, the considered pairs are dialog turns that are usually longer than just one sentence.

### v2

@@ Abstract
Best practices for high conflict conversations like counseling or customer support almost always include recommendations to paraphrase the previous speaker. Although paraphrase classification has received widespread attention in NLP, paraphrases are usually considered independent from context, and common models and datasets are not applicable to dialog settings. In this work, we investigate paraphrases across turns in dialog (e.g., Speaker 1: “That book is mine.” becomes Speaker 2: “That book is yours.”). We provide an operationalization of context-dependent paraphrases, and develop a training for crowd-workers to classify paraphrases in dialog. We introduce ContextDeP, a dataset with utterance pairs from NPR and CNN news interviews annotated for context-dependent paraphrases. To enable analysis on label variation, the dataset contains 5,581 annotations on 600 utterance pairs. We present promising results with in-context learning and with token classification models for automatic paraphrase detection in dialog.
@@ 1 Introduction
Dataset Acc. $\alpha$ Shortened Example Vote
@@
BALANCED 0.71 0.32 Guest: […] Maybe the money will help. Host: It can’t hurt, let’s put it that way. 9/20
@@
RANDOM 0.72 0.23 G: So both parties agree that we need to stop horrific acts of violence against animals. But everyone is standing behind this. It is time to stop horrific acts of brutality on animals. H: Britain’s Queen Elizabeth’s senior dresser writes "If her majesty is due to attend an engagement in particularly cold weather from 2019 onwards fake fur will be used to make sure she stays warm." it’s a very stark example of a monarch following public opinion in the U.K. which is moving away from fur and it very much embraces prevention of cruelty to the animals. 7/15
@@
PARA 0.65 0.19 G: […] it could be programmed in. But again, you’d have to set that up as part of your flight plan. H: So you’d have to say I’m going to drop to 5,000 feet, then go back up to 35,000 feet, and you would have had to have done that at the beginning. 8/15
@@
Table 1: Agreement Scores as an Indicator of Plausible Variation. For each dataset, we display the “accuracy” with the majority vote (Acc.) which is the mean overlap of a rater’s classification with the majority vote classification excluding the current rater and Krippendorff (1980)’s alpha ($\alpha$) for the binary classifications by all raters over all pairs. The relatively low K’s $\alpha$ scores can be explained by pairs where either label is plausible. We display such an example for each dataset with the share of annotators classifying it it as a paraphrase (Vote).
@@
Fortunately, in NLP, paraphrases have received wide-spread attention: Researchers have created numerous paraphrase datasets (Dolan and Brockett, 2005; Zhang et al., 2019; Dong et al., 2021; Kanerva et al., 2023), developed methods to automatically identify paraphrases (Zhang et al., 2019; Wei et al., 2022a; Zhou et al., 2022), and used paraphrase datasets to train semantic sentence representations (Reimers and Gurevych, 2019; Gao et al., 2021) and benchmark LLMs (Wang et al., 2018; bench authors, 2023). However, most previous work (1) has focused on context-independent paraphrases, i.e., texts that are semantically equivalent independent from the given context, and has not investigated the automatic detection of paraphrases across turns in dialog, (2) has classified paraphrases at the level of full texts even though paraphrases often only occur in portions of larger texts (see also Figure 1), (3) uses a small number of 1–3 annotations per paraphrase pair (Dolan and Brockett, 2005; Kanerva et al., 2023), (4) only annotate text pairs that are “likely” to include paraphrases using heuristics such as lexical similarity (Dolan and Brockett, 2005), although, especially for the dialog setting, we can not expect lexical similarity to be high for all or even most paraphrase pairs (e.g., the pair in Figure 1 only overlaps in two words) and (5) either use short annotation instructions Dolan and Brockett (2005) that rely on annotator intuitions or long and complex instructions Kanerva et al. (2023) that limit the total number of annotators.
@@
We address all five limitations with this work. First, we are, to the best of our knowledge, the first to focus on operationalizing, annotating and automatically detecting context-dependent paraphrases across turns in dialog. Dialog is a setting that is uniquely sensitive to context (Grice, 1957, 1975; Davis, 2002), e.g., “doing this for a while now” and “20th time […] as a television commentator” in Figure 1 are not generally semantically equivalent. Second, instead of classifying whether two complete texts A and B are paraphrases of each other, we focus on classifying whether there exists a selection of a text B that paraphrases a selection of a text A, and identifying the text spans that constitute the paraphrase pair (e.g., Figure 1). Third, we collect a larger number of annotations of up to 21 per item in line with typical efforts to address plausible human label variation (Nie et al., 2020; Sap et al., 2022). Even though context-dependent paraphrase identification in dialog might at first seem straight forward with a clear ground truth, similar to other “objective” tasks in NLP (Uma et al., 2021), human annotators (plausibly) disagree on labels (Dolan and Brockett, 2005; Kanerva et al., 2023). For example, consider the first text pair in Table 1. “[The money] can’t hurt” can be interpreted in at least two different ways: as a statement with approximately the same meaning as “the money will help” or as an opposing statement meaning the money actually won’t help but at least “It can’t hurt” either. Fourth, instead of using heuristics to select text pairs for annotations, we choose a dialog setting where paraphrases are relatively likely to occur: transcripts of NPR and CNN news interviews (Zhu et al., 2021) since in (news) interviews paraphrasing or more generally active listening is encouraged (Clayman and Heritage, 2002; Hight and Smyth, 2002; Sedorkin et al., 2023). While the interview domain shows some unique characteristics limiting generalizability (e.g., hosts using paraphrases to simplify the guest’s statements for the audience), the interview domain is is suitable to demonstrate our new task and includes a diverse set of topics and guests. Fifth, we develop an annotation procedure that goes beyond relying on intuitions and is scalable to a large number of annotators: an accessible example-centric, hands-on, 15-minute training before annotation.
@@
In short, we operationalize context-dependent paraphrases in dialog with a definition and an iteratively developed hands-on training for annotators. Then, annotators classify paraphrases and identify the spans of text that constitute the paraphrase. We release ContextDeP (Context-Dependent Paraphrases in news interviews), a dataset with 5,581 annotations on 600 utterance pairs from NPR and CNN news interviews. We use in-context learning (ICL) with generative models like Llama 2 or GPT-4 and fine-tune a DeBERTa token classifier to detect paraphrases in dialog. We reach promising results of F1 scores from $0.73$ to $0.81$. Generative models perform better at classification, while the token classifier provides text spans without parsing errors. We hope to advance dialog based evaluations of LLMs and the reliable detection of paraphrases in dialog. Code [^1] https://github.com/nlpsoc/Paraphrases-in-News-Interviews, annotated data [^2] https://huggingface.co/datasets/AnnaWegmann/Paraphrases-in-Interviews, [^3] This is in line with the license from the original data publication Zhu et al. (2021). and the trained model [^4] https://huggingface.co/AnnaWegmann/Highlight-Paraphrases-in-Dialog are publicly available for research purposes.
@@
Approximate Contextual Equivalence $\subseteq$ CP G: I’m like, "Fortnite", what is that? I don’t even know what it is – H: So, you weren’t even familiar?
@@ 2 Related Work
Paraphrase taxonomies commonly go beyond binary classifications to make more fine-grained distinctions between paraphrase types, often including considerations w.r.t. the context of the text pairs. Bhagat and Hovy (2013) and Kovatchev et al. (2018) describe substitutions and other lexical operations that result in paraphrases in a given sentential context. Shwartz and Dagan (2016) show that context information can reverse semantic relations between phrases. Vila et al. (2014) discuss text pairs that are equivalent when one presupposes encyclopedic or situational knowledge (e.g., referents or intentions555cases like ‘Close the door please” and “There is air flow”), but exclude them as non-paraphrases. Further, to the best of our knowledge, most previous work annotate sentence pairs without considering the document context, with Kanerva et al. (2023) being the only exception, and no previous work looking at detecting paraphrases in dialog.
@@ 3 Context-Dependent Paraphrases in Dialog
We define a context-dependent paraphrase as two text excerpts that are at least approximately equivalent in meaning in a given situation but not necessarily in all non-absurd situations.666definition combines elements from Kanerva et al. (2021) and Bhagat and Hovy (2013) For example, consider the first exchange in Table 2. In this situation, “I” uttered by the first speaker and “You” uttered by the second speaker are clearly signifying the same person. However, if uttered by the same speaker “I” and “you” probably do not signify the same person. The text pair in Table 2 is thus equivalent in at least one but not in all non-absurd situations. The text excerpts forming context-dependent paraphrases do not have to be complete utterances. In many cases they are portions of utterances, see highlights in Figure 1. Note that in dialog, the second speaker should rephrase part of the first speaker’s point in the given situation (context condition) and not just talk about something semantically related (equivalence condition).
@@
We specifically exclude common cases of disagreements between annotators777 derived from pilot studies, see also App. C.1 and specifically App. Table 15 that we consider not to be context-dependent paraphrases in dialog, see Table 3. First, we exclude text spans that might be considered approximately equivalent when they are looked at in isolation but do not represent a paraphrase of the guest’s point in the given situation (e.g., “the military” and “the army” in Table 3). Second, we exclude text pairs that diverge too much from the original meaning when the second speaker adds conclusions, inferences or new facts. In an interview setting, journalists make use of different question types and communication strategies relating to their agenda (Clayman and Heritage, 2002) that can sometimes seem like paraphrases. For example in Table 3, the host’s question “So, you …?” could be read as a paraphrase with the goal of checking understanding with the guest. However, it is more likely to be a declarative conclusion that goes beyond what the guest said.
@@ 4 Dataset
Generally, people do not paraphrase each other in every conversation. We focus on the news interview setting, because paraphrasing, or more generally active listening, is a common practice for journalists (Clayman and Heritage, 2002; Hight and Smyth, 2002; Sedorkin et al., 2023). We therefore also only consider whether the journalist (the interview host) paraphrases the interview guest and not the other way around. We use Zhu et al. (2021)’s MediaSum corpus which consists of over 450K news interview transcripts and their summaries from 1999–2019 NPR and 2000–2020 CNN interviews. [^8] Released for research purpose, see https://github.com/zcgzcgzcg1/MediaSum?tab=readme-ov-file.
@@ 4.2 Data Samples for Annotation
Even though paraphrases are relatively likely in the news interview setting, most randomly sampled text pairs still do not include paraphrases. To distribute annotation resources to text pairs that are likely to be paraphrase, previous work usually selects pairs based on heuristics like textual similarity features, e.g., word overlap, edit distance, or semantic similarity (Dolan and Brockett, 2005; Su and Yan, 2017; Dong et al., 2021). However, these approaches are systematically biased towards selecting more obvious, often lexically similar text pairs, possibly excluding many context-dependent paraphrases. For example, the guest and host utterance in Figure 1 have varying lengths, only overlap in three words and have a semantic similarity score of only 0.13999 using cosine-similarity and encodings from https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Similar to Kanerva et al. (2023), we instead use a manual selection of promising text pairs for annotation: We (1) randomly sample a set of text pairs and (2) manually classify at each of them to (3) select three sets of text pairs that vary in their paraphrase distribution for the more resource-intensive crowd-sourced annotations: the RANDOM, BALANCED and PARA set.
@@
Lead Author Annotation. We shuffle and uniformly sample 1,304 interviews. For each interview, we sample a maximum of 5 consecutive (guest, host)-pairs. To select promising paraphrase candidates, the lead author then manually classifies all 4,450 text pairs as paraphrases vs. non-paraphrases (see App. B.2 for details).101010 After experimenting with crowd-workers, having a first pass for selection done by one of our team seemed the best considering cost-performance trade-offs. In total, about 14.9% of the sampled text pairs are classified as paraphrases by the lead author. On a random set of 100 (guest, host)-pairs (RANDOM), we later compare the lead author’s classifications with the crowd-sourced paraphrase classifications (see App. 13). 89% of the lead author’s classifications are the same as the crowd majority. Note that the lead author’s classifications do not affect the quality of the annotations released with the dataset but only the text pairs that are selected for annotation. However, using lead author annotations instead of lexical level heuristics should increase paraphrase diversity in the released dataset beyond high lexical similarity pairs.
Split # (guest, host)-pairs # annotations
Train 420 3896
Dev 88 842
Test 92 843
Total 600 5,581
@@ 5.1 Annotation Task
Given a (guest, host) utterance pair, annotators (1) classify whether the host is paraphrasing any part of the guest’s utterance and, if so, (2) highlight the paraphrase in the guest and host utterance. This results in data points like the one in Figure 1. Note that our setup differs from prior work, which usually involves classifying whether an entire text B is a paraphrase of an entire text A (e.g., Dolan and Brockett, 2005). Instead, given texts A and B, our task is to determine whether there exists a selection of words from text B and text A, where the selection of text B is a paraphrase of the selection of text A. Our annotators are not only performing binary classification, but they also highlight the position of the paraphrase. To the best of our knowledge, we are the first to approach paraphrase detection in this way. Moreover, in contrast to previous work, the considered text pairs are usually longer than just one sentence and are contextualized dialog turns.
5.2 Plausible Label Variation
The task of annotating context-independent paraphrases is already difficult. Disagreements between human annotators are common (Dolan and Brockett, 2005; Krishna et al., 2020; Kanerva et al., 2023) — even with extensive manuals for annotators (Kanerva et al., 2023). In related semantic tasks like textual entailment,111111 Paraphrase classification has been repeatedly equated to (bi-)directional entailment classification (Dolan and Brockett, 2005; Androutsopoulos and Malakasiotis, 2010) disagreements have been linked to plausible label variations inherent to the task (Pavlick and Kwiatkowski, 2019; Nie et al., 2020; Jiang and de Marneffe, 2022).